OpenStudy (anonymous):

PLEASE HELP!!!! A customer in a restaurant wants to leave a tip using a penny, nickel, dime, quarter, and a half dollar. If the customer insists on leaving only two coins, how many different possibilities are there?

OpenStudy (anonymous):

by adding the $ amount up and then divide by 2

OpenStudy (anonymous):

which will be 45.5 ????

OpenStudy (anonymous):

p+n+d+q+Half dollar ------------------- 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

no problem, anymore questions

OpenStudy (anonymous):

feel free to ask me anymore questions

OpenStudy (anonymous):

ok i have one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

From a bag containing 6 white balls, 3 black balls and 13 red balls , two balls are drawn at random. What is the probability that they are both white?

OpenStudy (anonymous):

1) .078 2) .074 3) .065 4) .062

OpenStudy (anonymous):

add up all ball numbers and then put 2 over 6+3+13=? 2/?

OpenStudy (anonymous):

.065?

OpenStudy (anonymous):

no, try again

OpenStudy (anonymous):

6+3+13= what

OpenStudy (anonymous):

22

OpenStudy (anonymous):

yes, so put 22 under 2 and your answer will be 2/22= 2 -- 22

OpenStudy (anonymous):

you may have to simplify it u may not